Jefferson Steel requires $15 million to fund its current year’s capital projects. Jefferson will finance part of its needs with equity. The firm’s common stock is selling in the market at $180 per share. Dividends of $1 per share were recently paid (Do). Dividend growth of 6 percent per year is expected for the foreseeable future. The market is currently demanding a 6 percent premium on the average risk stock and Tbonds are currently yielding 3%. Preferred stock is selling in the market at $97 per share. The preferred has a dividend rate of 4.5% percent on $200 face value per share. The remainder of needs will be financed with debt. Jefferson's current bonds are ten year, $1000 par bonds with a coupon rate of 8% and are selling at a price of $1020 per bond. Interest is paid semi-annually on the bonds. The firm faces a 30 percent marginal tax rate and the project asset beta is estimated at 1.45. Jefferson believes the capital asset pricing model is the best way to estimate the cost of equity. Jefferson’s balance sheet, in part, is presented below. The number of units outstanding are shown in parenthesis.

($ x millions) Accounts Payable Accruals LT Debt ($1000 par) Total Debt $17,000,000 Pfd Stock 14,000,000 Common stock @ par (1. Determine the weights of debt, preferred stock and common equity to be used in financing the project.

Wd =

Wp =

We =

2. Determine the cost of capital to the firm:

WACC =
Answer #1

1. total value of capital from balance sheet = Stock + debt+ preferred stock = 8250000+15000000+25000000 = 48250000

weight of debt , wd= 25000000/48250000 = 0.17

weight of preferred stock = 0.52

weight of equity = 0.31

2. wacc = we*re + wd*rd + wp*rp = 0.055740

given

rd 8%
re 0.0735
rp 4.50%
t 30%